OpenStudy (anonymous):

hi, for the expansion of power series (logarithmitic series) , ln(1+X) , why the condition for x is between -1 and 1 which x can be 1 but x cant be -1 ? 2. Relevant equations

OpenStudy (aravindg):

Because log is not defined for negative numbers?

OpenStudy (aravindg):

oh wait ..I read it wrong. 1 sec.

OpenStudy (yttrium):

Can you revise the question?

zepdrix (zepdrix):

`hi, for the expansion of power series (logarithmitic series) , ln(1+X)` Do you mean ln(1-x) as is pictured?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Mmm lemme just write some stuff out and see if I can make sense of this.

zepdrix (zepdrix):

Since, \[\Large\bf\sf \frac{d}{dx}\ln(1-x)=\frac{-1}{1-x}\]That implies that,\[\Large\bf\sf \ln(1-x)=\int\limits \frac{-1}{1-x}\;dx\quad=\quad -\int\frac{1}{1-x}\;dx\]And then we can write that as a geometric series,\[\Large\bf\sf =-\int\limits \sum_{n=0}^{\infty}x^n\;dx\]

OpenStudy (anonymous):

thanks, but wat do u mean? can u try to type out using LA tex ?

zepdrix (zepdrix):

\[\Large\bf\sf =-\sum_{n=0}^{\infty} \int\limits x^n\;dx\quad=\quad -\sum_{n=0}^{\infty}\frac{x^{n+1}}{n+1}\]

zepdrix (zepdrix):

So then those terms are giving us,\[\Large\bf\sf -x-\frac{x^2}{2}-\frac{x^3}{3}-...\]And the part we're confused on is the radius of convergence or something?

OpenStudy (anonymous):

i can only see a bunch of equation... wat does large and bf , sf-x means?

zepdrix (zepdrix):

Are you using Internet Explorer? You probably should switch browsers. It doesn't support the LaTeX plugin very well. The equations aren't showing up correctly on your screen.

OpenStudy (anonymous):

http://www.codecogs.com/latex/eqneditor.php i men this one...

OpenStudy (anonymous):

i'm using chrome right now.....

zepdrix (zepdrix):

We have the mathjax plugin on this site. The equations should show up correctly on here. I'm not sure why it's not working for you..

OpenStudy (anonymous):

mathjax plug in? need to download? how to use it?

zepdrix (zepdrix):

No it should work automatically :( You're on chrome? Hmm that's weird..

OpenStudy (anonymous):

can you suggest what should i do now to view the equation?

zepdrix (zepdrix):

Is this your first time using the site? Maybe you just need a reboot or something +_+ Blah I dunno!! This is an old old thread but matt's response(tips for fixing latex display) near the top might still be relevant. http://openstudy.com/study#/updates/4f4aa4c4e4b00c3c5d337cd2

OpenStudy (anonymous):

thanks ! i can view it now after rebooting... i will view it later... currently busy with my other homework...

zepdrix (zepdrix):

So I wrote it as a geometric series. Recall that the geometric series only converges when that term is less than 1.\[\Large\bf\sf |x|\lt1\]Which we can write as,\[\Large\bf\sf -1\lt x\lt 1\]Hmm but that still doesn't explain the allowance for the lower bound.. hmmm

OpenStudy (anonymous):

Maybe I misunderstand this question, but the way I read it all you're asking is to validate the regions of convergence, which is a trivial thing to do, but needs a careful analysis of the so called upper and lower bounds, or region endings. \[\large \log(1+x) = \sum_{n=1}^{+ \infty} \frac{(-1)^{n+1}}{n}x^n \] so for \(x=1\) (only focus on the right hand side) we get \[ \large \sum_{n=1}^{+ \infty} \frac{(-1)^{n+1}}{n}\] which converges, it's the alternating harmonic series, and for \(x=-1\) we get \[\large \sum_{n=1}^{+ \infty} \frac{(-1)^{n+1}(-1)^n}{n}= - \sum_{n=1}^{+ \infty} \frac{1}{n} \] which diverges, because it's the harominc series.

OpenStudy (anonymous):

If your question concerns the shifted \( \log(1-x)\) then you should already, from my example above, be able to see why it now works the other way around, meaning that it converges for \( -1 \leq x <1 \)

OpenStudy (anonymous):

if x=1 then shoudnt the RHS =1^(n) , why you left it out?

OpenStudy (anonymous):

\(1^n=1, \forall n \in \mathbb{Z}\), we sum over \(\mathbb{N}\) so you can say that: \[1^1=1^2=1^3= \dots =1^n = 1 \]

OpenStudy (anonymous):

You can't include -1 because \(\ln(-1+1)=\ln0\) is undefined.